OpenStudy (anonymous):

If y varies directly with x, write an equation for the direct variation. If y=6 when x=24, what is the value of x when y=7? Is the answer Y=56?

OpenStudy (0487308):

28

OpenStudy (anonymous):

Can you explain how you got the answer please?

OpenStudy (0487308):

y = \[\frac{ x }{ 4 }\]

OpenStudy (0487308):

This means that 7 = x/4. Multiply each side my 4, and you get 28.

OpenStudy (anonymous):

Side?

OpenStudy (0487308):

Side of the equation, with the center of the equation being the equal sign.

OpenStudy (0487308):

Do you get it now?

OpenStudy (anonymous):

Ya....

OpenStudy (anonymous):

@mathmale is 28 right?

OpenStudy (0487308):

Who's @mathmale

OpenStudy (mathmale):

"If y=6 when x=24," write the direct proportion that applies here. In other words, find k in y=kx. Once you've done that, I can quickly check your work by substituting y=7. Better yet, start with y our own y=kx and substitute y=7. What is x?

OpenStudy (0487308):

@mathmale Aren't there infinitely many possible equations to represent this, because there is only one ordered pair given? In order for you to have one absolute equation to represent a graph, you need two or more given ordered pairs (supposing that the equation is linear).

OpenStudy (anonymous):

@mathamale Is this correct? y=kx 7=ktimes8 k=8 y=8x y=7times8 y=56

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

Is the equation correct?

OpenStudy (anonymous):

@Mehek14 Is this equation correct? Please scroll to the bottom! :D

OpenStudy (anonymous):

Is the answer y=56?

OpenStudy (anonymous):

THANK YOU! :D

TheSmartOne (thesmartone):

I believe you have the wrong answer.

TheSmartOne (thesmartone):

I don't know what Mehek is saying, but your answer is absolutely wrong.

OpenStudy (anonymous):

? @TheSmartOne

OpenStudy (anonymous):

How is it wrong @TheSmartOne

TheSmartOne (thesmartone):

It's all wrong.

TheSmartOne (thesmartone):

All your work, all the numbers you used. They are all just wrong.

OpenStudy (anonymous):

Ok no need to rub it in...But can you explain?

TheSmartOne (thesmartone):

You need to learn how to deal with specific information first, before jumping to what you need to find. "If y varies directly with x, write an equation for the direct variation. If y=6 when x=24" Now, many people have posted this above, but can you tell me the equation for direct variation?

OpenStudy (anonymous):

y=kx

TheSmartOne (thesmartone):

ok, now here's the first part of your question: "If y varies directly with x, write an equation for the direct variation. If y=6 when x=24" Can you substitute y = 6, and x = 24 into it and solve for k?

TheSmartOne (thesmartone):

The silence is real...

OpenStudy (anonymous):

Ummmm...Can you tell me what you mean by substitute? @TheSmartOne

TheSmartOne (thesmartone):

You're doing it all wrong @mathmale -.-

OpenStudy (anonymous):

Im so confused ^_^

TheSmartOne (thesmartone):

@mathmale You substitued the wrong numbers in and got k = 24/7 If you properly substituted it, you would have gotten 6/24 = 1/4

OpenStudy (mathmale):

Borrowing TheSmartOne's input: ok, now here's the first part of your question: "If y varies directly with x, write an equation for the direct variation. If y=6 when x=24" Can you substitute y = 6, and x = 24 into it and solve for k? Direct variation implies that y=kx. If y = 6 when x=24, then 6=k(24), and k=(6/24), or 1/4. Thus, y = (1/4) x. Check! If x=24, is y=6 as it should be?

OpenStudy (mathmale):

Yes or no?

OpenStudy (mathmale):

Assuming that our equation of direct variation is y=(1/4)x, and now letting y=7, find x. Just substitute 7 for y in y= (1/4) x. x=?

OpenStudy (anonymous):

I still can't tell whose correct

TheSmartOne (thesmartone):

we're all on the same page now :)

OpenStudy (mathmale):

@Angel360 : I went through the calculation of the value of k and then checked my own work. Next, I let y take on a different value (7 instead of 6) and substituted that into my direct variation equation. If y = (1/4)x, and y=7, find x. Please demonstrate that you can do this.